# Pushing the Limits of Distillation-Based Class-Incremental Learning via Lightweight Plugins

### v2

@@ 4.1 Task-Specific Plugins Deployment
In DLC, the plugins and $\phi(\boldsymbol{x})$ undergo a decoupled two-phase training procedure. Specifically, when a new task $t$ arrives, we first train the base model following the underlying baseline procedure, updating only $\phi(\boldsymbol{x})$ and the classifier $\boldsymbol{W}$ while keeping all plugins frozen. After this phase, $\phi(\boldsymbol{x})$ is frozen to prevent knowledge drift, and $L_{t}$ is trained with the same data. Upon completion, $L_{t}$ is permanently frozen to safeguard the acquired task-specific knowledge. The training objective for all plugins incorporates the cross-entropy loss and an auxiliary loss [33] commonly used in expansion-style methods, formulated as:
@@
The two-phase procedure ensures that plugins are trained without interference from distillation losses, enabling specialized learning for the target task. Since $\phi(\boldsymbol{x})$ remains frozen during plugin updates, plugin training is stable, and it also prevents the baseline training of $\phi(\boldsymbol{x})$ from being affected by plugin training, preserving the plug-and-play property of DLC.
During inference, all plugins are activated to extract their respective feature representations. These representations are then concatenated to form an integrated feature vector, which enhances the base model through a multi-task feature ensemble. For a linear classifier $\boldsymbol{W}$, the classification process can be formulated as:
$\hat{y}=\boldsymbol{W}[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),...,\phi^{L_{T}}(\boldsymbol{x})]$ (6)
@@
where $T$ denotes the maximum index of known tasks. Notably, we do not need to run the full extractor $T$ times: each plugin injects a residual only at its deployment layer, so the backbone activations up to the earliest deployed layer are shared and computed once, avoiding redundant computation and improving inference efficiency.
4.2 LoRA-Style Residual Plugin
@@
To instantiate each task-specific plugin with minimal parameters, we adopt a LoRA [13]-style low-rank residual parameterization. This design is well-suited for plugin deployment: the rank $r$ provides a simple capacity knob under a fixed budget, the same low-rank residual form applies to both linear and convolutional operators, and the residual structure supports our decoupled training and inference aggregation since plugins can be independently trained, frozen, and combined without modifying the backbone. We emphasize that the backbone weights are kept intact throughout, and the low-rank component is introduced as an auxiliary module that adds a task-specific residual to the feature. That is, the LoRA plugin does not replace or factorize the original weights, but injects additional task knowledge via an additive low-rank branch, which aligns with our plug-and-play deployment setting.
@@
Since most CIL pipelines train CNN-based backbones from scratch, we instantiate the plugin for convolutional operators using a LoRA-style low-rank residual branch (implemented as two convolutions). For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation. The LoRA plugin first processes the input with $\mathbf{A}$:
@@
where $\mathbf{B}$ $\in\mathbb{R}^{C{\text{out}}\times r\times 1\times 1}$ restores the channel dimensionality, and $\boldsymbol{h}_{\text{lora}}$ is added as a residual to the layer output.
In summary, the forward pass of a convolutional layer equipped with the LoRA-style residual plugin is:
@@
where $\alpha$ is a scaling hyperparameter, and $r$ is the rank.
Compared to the model expansion, instantiating task-specific plugins with a LoRA-style parameterization can bring significant parameter advantages. Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while the plugin only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$. For expanding one layer of CNN over $T$ tasks, considering that in a CNN in CIL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7). Since $r$ is low-rank, it can also be regarded as a constant (generally set to 8 or 16). Then, using LoRA-style plugin instantiation reduces the parameter-space complexity of expanding a single CNN layer from $\Theta(TC^{2})$ to $\Theta(TC)$. Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can substantially reduce the number of parameters.
@@ 4.3 Dynamic Weighting Unit
For a sample from task $t$, all the plugins trained on other tasks are considered non-target plugins. Given that the replay buffer contains only a minimal fraction of the original data for each task, these non-target plugins are not trained on the full data from task $t$. In particular, plugins trained on tasks 1 to t1t-1 never encounter samples from the task $t$. Consequently, the residuals extracted by these plugins could be largely irrelevant to the task $t$, thus introducing uninformative features or even detrimental noise to the classifier.
To mitigate the influence of uninformative residuals from non-target plugins, we will create a weighting unit corresponding to concentrated representations when training on a new task. This unit comprises two linear layers followed by a sigmoid function $\sigma(\cdot)$, denoted as $\omega(\mathbf{x})=\sigma\Big(\mathbf{W}_{2}\text{ReLU}\big(\mathbf{W}_{1}\mathbf{x}\big)\Big),\mathbf{x}\in\mathbb{R}^{k},\mathbf{W}_{1}\in\mathbb{R}^{k\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times k}$, where the sigmoid function normalizes the output to the [0, 1] range, indicating their importance weights. The concatenated representations are passed through this unit before being fed into the classifier, then weighted representations is derived from $\tilde{\mathbf{h}}^{(i)}=\mathbf{\omega}^{(i)}\odot\mathbf{h}^{(i)}$, where $\odot$ is element-wise multiplication operator. The weighting unit can be regarded as a gate for the classifier. Jointly trained with the classifier, it helps the model adapt to the high-dimensional feature space and automatically assigns lower weights to uninformative or noisy features.
@@
The automatically trained weighting unit fails to explicitly model the varying importance of representations from different tasks. For a sample $i$ in task $t$, we denote the representations adapted by LoRA plugins trained on tasks 1 to t1t-1 as $\mathbf{h}^{(i)}_{\text{pre}}$ and those adapted by the remaining LoRA plugins as $\mathbf{h}^{(i)}_{\text{pos}}$, respectively. Then $\mathbf{h}^{(i)}_{\text{pre}}$ should be assigned lower weights as corresponding LoRA plugins were not optimized on task $t$. However, since the ground-truth task index of each sample is unavailable during the CIL testing, we cannot explicitly control the weighting unit based on task identity. Instead, we employ a supervised regression training to guide the weighting unit to approximate the desired weighting scheme. We decompose the weight $\mathbf{\omega}^{(i)}$ for sample $i$ into $\mathbf{\omega}^{(i)}_{\text{pre}},\mathbf{\omega}^{(i)}_{\text{pos}}$] Then the decision process of the linear classifier can be rewritten as follows:
$\boldsymbol{q}^{i}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})+\mathbf{W}_{\text{pos}}(\mathbf{\omega}^{(i)}_{\text{pos}}\odot\mathbf{h}^{(i)}_{\text{pos}})$. (10)
@@
Formally, the contribution of $\mathbf{h}^{(i)}_{\text{pre}}$ to the classifier’s logits can be expressed as $\mathbf{\xi}=\mathbf{W}_{\text{pre}}(\mathbf{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})$. To generate effective weights, an importance-aware optimization objective could be incorporated into the training process, denoted as $\mathcal{L}_{\text{IA}}$. This loss function employs the Mean Squared Error (MSE), expressed as follows:
$\mathcal{L}_{\text{IA}}=\frac{1}{k}{\|\mathbf{\omega}-\mathbf{\omega}_{\text{ideal}}\|}_{2},$ (11)
@@
where the $\mathbf{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\mathbf{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\mathbf{\omega}^{(i)}_{\text{ideal-pre}}=0$. In the ideal training case, the model has $\mathbf{\omega}^{(i)}_{\text{pos}}\to 1$ and $\mathbf{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits. Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed. After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the plugin training, and $\mathcal{L}_{\text{plg}}$ in Eq. (5) is extended as follows:
@@
Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison. ‘-DLC’ indicates the method using the DLC framework. ‘Extra/Feat #P’ denotes the total number of extra parameters required for deploying DLC and the feature extractor, respectively.
@@ 5 Experiments
This section conducts extensive experiments. Sec. 5.2 evaluates the validity of DLC on distillation-based baselines. Sec. 5.3 investigates the performance of SOTA CIL methods under the memory-aligned case. Sec. 5.4 evaluates the compatibility of DLC with other data augmentation and plug-and-play loss functions for CIL. Sec. 5.5 analyzes the additional inference overhead introduced by DLC in terms of FLOPs and latency.
@@ 5.1 Experimental Setup
Datasets and Split. We chose three benchmark datasets commonly used in CIL, including CIFAR-100 [15], ImageNet-100 [9], and Tiny-ImageNet-200 [28], where the number following the dataset is the total number of classes included. Following the ‘B/Base-m, Inc-n’ rule proposed by [7], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 Inc5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first incremental task, and $n$ represents that of every remaining stage. For a fair comparison, we ensure that the training and test datasets are identical for all methods. Detailed information about the dataset is placed in the Appendix section.
@@
Implementation Details. All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [41], and Gao et al. [10]. We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets. The optimizer employs SGD with a fixed learning rate of 0.1, incorporating a multi-step schedule. The detailed training hyperparameters, like the optimizer and batch size, largely follow the default settings in PyCIL. All the methods are run using the same random number seed. To eliminate potential confounding effects from increased training epochs, all baselines are run under both the PyCIL-recommended epoch setting and the total epoch budget of DLC’s two-stage training, and the maximum accuracy is reported.
@@
Evaluation Metrics. Following the benchmark protocol settings [23], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ tasks, and $A_{T}$ is the last accuracy on the test set that includes all learned tasks. We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
5.2 Effectiveness on Distillation-Based CIL

### v3

@@ 4.1 Classifier-Proximal Plugin Deployment
The plugins and $\phi(\boldsymbol{x})$ are trained in a decoupled two-phase manner. When a new task $D_{t}$ arrives, we first follow the underlying baseline to train the backbone and classifier while keeping all plugins frozen. After this phase, the backbone is frozen, and only the newly introduced plugin $L_{t}$ is optimized on the same data. Once trained, $L_{t}$ is permanently frozen to preserve the acquired stage-specific knowledge. The plugin objective combines the cross-entropy loss with an auxiliary loss commonly adopted in expansion-style methods [38]:
@@
This decoupled procedure allows plugins to learn stage-specific corrections without being directly entangled with distillation optimization, while also preventing plugin training from interfering with the baseline update of the backbone.
During inference, all learned plugins are activated to produce their corresponding plugin-enhanced representations. Because each plugin encodes a different stage-specific residual refinement on top of the shared backbone output, these representations provide diverse semantic cues with potentially complementary information. Their aggregation can thus be viewed as a classifier-proximal form of multi-stage feature reuse, in spirit similar to DenseNet [15] feature reuse, enriching the final prediction with knowledge accumulated across learning stages. For a linear classifier $\boldsymbol{W}$, the classification process can be written as:
$\hat{y}=\boldsymbol{W}[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{T}}(\boldsymbol{x})],$ (6)
@@
where $T$ denotes the total number of learned stages so far. Notably, the full feature extractor does not need to be executed $T$ times: since each plugin injects a residual after the last layer, the backbone activations before the deployed layer can be shared and computed once, avoiding redundant computation.
For continual learning to remain practical in real-world scenarios, the goal is not to allocate an extractor-level or layer-level plugin for every learning stage, but to introduce a lightweight auxiliary module plugin with minimal storage overhead. Such a design naturally calls for parameter-efficient fine-tuning (PEFT) modules [11]. However, these modules have previously been studied mostly in transfer learning, where they are typically deployed on top of a knowledge-rich pretrained backbone. Distillation-based continual learning presents a fundamentally different regime: the backbone is not a static pretrained model with abundant prior knowledge, but a continually updated learner whose knowledge is progressively accumulated and preserved through replay and distillation. As discussed in Sec. 3.2, knowledge distillation together with a replay buffer helps retain previously acquired knowledge, thereby constraining distribution drift during continual updates. This makes it possible to revisit PEFT modules in a dynamic backbone. Concretely, after training on stage $t$, the output deviation at layer $\ell$ is bounded by $K_{\ell}\Gamma_{t}$, where $K_{\ell}$ depends on hyperparameters such as the network architecture and distillation temperature, and $\Gamma_{t}$ reflects the distribution discrepancy induced by the adopted replay and distillation strategy. This bound directly limits the worst-case feature drift on the old-stage distribution, enabling stable optimization of PEFT modules in this dynamic regime. A detailed proof is provided in the supplementary material.
4.2 LoRA-Style Plugins
@@
To instantiate each stage-specific plugin with minimal parameters, we adopt a LoRA [14]-style low-rank residual parameterization. This design is well-suited for plugin deployment: the residual structure supports our decoupled training and inference aggregation since plugins can be independently trained, frozen, and combined without modifying the backbone. We emphasize that the backbone weights are kept intact throughout, and the low-rank component is introduced as an auxiliary module that adds residual to the feature. That is, the LoRA plugin does not replace or factorize the original weights, but injects additional knowledge via an additive low-rank branch, which aligns with our plug-and-play deployment setting.
@@
Following the classifier-proximal deployment strategy in Sec. 4.1, this plugin is attached to the final feature layer of $\phi(\boldsymbol{x})$; for CNN backbones, it is instantiated on the representation produced by the last convolutional layer. However, LoRA was originally proposed for Transformer models, while distillation-based CL methods are often built upon CNN-based backbones trained from scratch. We therefore provide a convolutional LoRA-style plugin instantiation for this setting, implemented as a lightweight low-rank residual branch with two convolutions. For a convolutional layer with weight tensor $W_{\text{conv}}\in\mathbb{R}^{C_{\text{out}}\times C_{\text{in}}\times K\times K}$ and input $\boldsymbol{x}\in\mathbb{R}^{N\times C_{\text{in}}\times H\times W}$, the original forward pass is $\boldsymbol{h}=W_{\text{conv}}\ast\boldsymbol{x}$, where $\ast$ denotes the convolution operation. The LoRA plugin first processes the input with $\mathbf{A}$:
@@
where $\mathbf{B}$ $\in\mathbb{R}^{C{\text{out}}\times r\times 1\times 1}$ restores the channel dimensionality, and $\boldsymbol{h}_{\text{lora}}$ is added as a residual to the layer output. In summary, the forward pass of a convolutional layer equipped with the LoRA-style residual plugin is:
@@
where $\alpha$ is a scaling hyperparameter, and $r$ is the rank. Following [12], $r$ and $\alpha$ are set to the recommended values of 8 and 16.
Compared to the model or layer-level expansion, instantiating the LoRA-style PEFT plugins can bring significant parameter advantages. Specifically, for a convolutional layer ($C_{in}$, $C_{out}$, $K\times K$), the space complexity of its parameters is $\Theta(C_{in}C_{out}K^{2})$, while the plugin only occupies $\Theta(r(C_{in}K^{2}+C_{out}))$. For expanding one layer of CNN over $T$ stages, considering that in a CNN in CL, $C_{out}$ is generally greater than or equal to $C_{in}$, we set $C_{in}=C_{out}=C$, and $K$ can be regarded as a constant (usually 3 or 7). Moreover, since the low rank $r$ is set to 8, it can be treated as a constant. Then, using LoRA-style plugin instantiation reduces the parameter-space complexity of expanding a single CNN layer from $\Theta(TC^{2})$ to $\Theta(TC)$. Considering that $C$ is generally between 64 and 512, reducing its order of magnitude from quadratic to linear can substantially reduce the number of parameters.
@@ 4.3 Dynamic Weighting Unit
For any sample from stage $t$, all the plugins trained on other tasks $D_{i\neq t}$ are considered non-target plugins. Given that the replay buffer contains only a minimal fraction of the original data, these non-target plugins are not trained on the full $D_{t}$. In particular, plugins trained on stages 1 to t1t-1 never encounter samples from $D_{t}$. Consequently, the residuals extracted by these plugins could be largely irrelevant to the stage $t$, thus introducing uninformative features or even detrimental noise to the classifier. To mitigate their influence, we will create a weighting unit corresponding to concentrated representations when training on a new stage. This unit comprises two linear layers followed by a sigmoid function $\sigma(\cdot)$, denoted as $\omega(\mathbf{x})=\sigma\Big(\mathbf{W}_{2}\text{ReLU}\big(\mathbf{W}_{1}\mathbf{x}\big)\Big),\mathbf{x}\in\mathbb{R}^{k},\mathbf{W}_{1}\in\mathbb{R}^{k\times d},\mathbf{W}_{2}\in\mathbb{R}^{d\times k}$, where $\sigma(\cdot)$ normalizes the output to the [0, 1] range. During inference, we first compute the importance weights for the aggregated plugin-enhanced representation:
$\boldsymbol{\omega}=\omega\!\left([\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{T}}(\boldsymbol{x})]\right),$ (10)
then perform inference using the reweighted representation:
$\hat{y}=\boldsymbol{W}\left(\boldsymbol{\omega}\odot[\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{T}}(\boldsymbol{x})]\right).$ (11)
Where the $\odot$ denotes element-wise multiplication. The weighting unit can be regarded as a gate for the classifier. Jointly trained with the classifier, it helps the model adapt to the high-dimensional feature space and automatically assigns lower weights to uninformative or noisy features.
@@
The automatically trained weighting unit fails to explicitly model the varying importance of representations from different stages. For a sample $i$ in stage $t$, we denote the representations adapted by plugins trained on stages 1 to t1t-1 ($\phi^{L_{1}}(\boldsymbol{x}),\phi^{L_{2}}(\boldsymbol{x}),\ldots,\phi^{L_{t-1}}(\boldsymbol{x})]$) as $\mathbf{h}^{(i)}_{\text{pre}}$ and those adapted by the remaining plugins as $\mathbf{h}^{(i)}_{\text{pos}}$, respectively. Then $\mathbf{h}^{(i)}_{\text{pre}}$ should be assigned lower weights as corresponding plugins were not optimized with samples in $D_{t}$. Since the stage of each sample is unavailable during the CL testing, we cannot explicitly control the weighting unit based on stage identity. Instead, we employ a supervised regression training to guide the weighting unit to approximate the desired weighting scheme. We decompose the weight $\boldsymbol{\omega}^{(i)}$ for sample $i$ into $\boldsymbol{\omega}^{(i)}_{\text{pre}},\boldsymbol{\omega}^{(i)}_{\text{pos}}$] Then the decision process of the linear classifier can be rewritten as follows:
$\boldsymbol{q}^{i}=\mathbf{W}_{\text{pre}}(\boldsymbol{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})+\mathbf{W}_{\text{pos}}(\boldsymbol{\omega}^{(i)}_{\text{pos}}\odot\mathbf{h}^{(i)}_{\text{pos}})$. (12)
@@
Formally, the contribution of $\mathbf{h}^{(i)}_{\text{pre}}$ to the classifier’s logits can be expressed as $\mathbf{\xi}=\mathbf{W}_{\text{pre}}(\boldsymbol{\omega}^{(i)}_{\text{pre}}\odot\mathbf{h}^{(i)}_{\text{pre}})$. To generate effective weights, an importance-aware optimization objective could be incorporated into the training process, denoted as $\mathcal{L}_{\text{IA}}$. This loss function employs the Mean Squared Error (MSE), expressed as follows:
$\mathcal{L}_{\text{IA}}=\frac{1}{k}{\|\boldsymbol{\omega}-\boldsymbol{\omega}_{\text{ideal}}\|}_{2},$ (13)
@@
where the $\boldsymbol{\omega}_{\text{ideal}}$ is the ideal weight output, it has $\boldsymbol{\omega}^{(i)}_{\text{ideal-pos}}=1$ and $\boldsymbol{\omega}^{(i)}_{\text{ideal-pre}}=0$. In the ideal training case, the model has $\boldsymbol{\omega}^{(i)}_{\text{pos}}\to 1$ and $\boldsymbol{\omega}^{(i)}_{\text{pre}}\to 0$, then it has $\mathbb{E}[\mathbf{\xi}]\to 0$ and $\text{Var}[\mathbf{\xi}]\to 0$, so the $\mathbf{h}^{(i)}_{\text{pre}}$ have negligible influence on prediction logits. Even if the ideal scenario is hard to achieve, the classifier can still produce more accurate predictions as long as the representations adapted by reliable plugins are assigned higher weights, while those adapted by non-target plugins are appropriately suppressed. After enabling the weighting unit, $\mathcal{L}_{\text{IA}}$ is incorporated into the plugin training, and $\mathcal{L}_{\text{plg}}$ in Eq. (5) is extended as follows:
@@
Table 1: Last performance $A_{T}$ and average performance $\overline{A}$ comparison.
@@
Figure 3: Comparison of $A_{T}$ with/without DLC when enhanced with AutoAugment (AA) , $\mathcal{L}_{\text{inter-intra}}$ and C-flat.
@@ 5 Experiments
This section conducts extensive experiments. Sec. 5.2 evaluates the validity of DLC on distillation-based baselines. Sec. 5.3 further evaluates DLC under challenging CL settings, including long-sequence, few-shot, and class-imbalanced scenarios. Sec. 5.4 investigates the performance of SOTA CL methods under the memory-aligned case. Sec. 5.5 analyzes the additional inference overhead introduced by DLC in terms of FLOPs and latency.
@@ 5.1 Experimental Setup
Datasets and Split. We chose three benchmark datasets commonly used in CL, including CIFAR-100 [17], Tiny-ImageNet-200 [32], and the large-scale ImageNet-100 [7]. Following the ‘B/Base-m, Inc-n’ rule proposed by [5], we slice each of the above three datasets into CIFAR-100 B10 Inc10, CIFAR-100 B50 Inc10, CIFAR-100 B5 Inc5, ImageNet-100 B50 Inc10, and Tiny-ImageNet B40 Inc40 for the experiment, where $m$ is the number of classes in the first learning stage, and $n$ represents that of every remaining stage. For a fair comparison, we ensure that the training and test datasets are identical for all methods. Detailed information about the dataset is provided in the Appendix.
@@
Implementation Details. All methods are implemented in PyTorch, with the baseline methods referencing the PyCIL [46], and Gao et al. [8]. We employ the lightweight ResNet-32 tailored for CIFAR datasets on CIFAR-100, while utilizing ResNet-18 for all other benchmark datasets. The optimizer employs SGD incorporating the multi-step schedule, with a learning rate of 0.1. The detailed training hyperparameters, like the optimizer and batch size, largely follow the default settings in PyCIL. All the methods are run using the same random seed. To eliminate potential confounding effects from increased training epochs, all baselines are run under both the PyCIL-recommended epoch setting and the total epoch budget of DLC’s two-stage training, and the maximum accuracy is reported.
@@
Evaluation Metrics. Following the benchmark protocol settings [27], we use $A_{b}$ to denote the $b$-stage accuracy on the test set that includes all known classes after trained with $D_{1},D_{2},\cdots,D_{b}$, $\overline{A}=\frac{1}{T}\sum_{b=1}^{T}A_{b}$ is average accuracy over $T$ stages, and $A_{T}$ is the last accuracy on the test set that includes all learned stages. We use $\overline{A}$ and $A_{T}$ to evaluate the model’s performance, which reflects the model’s dynamic learning capacity and generalization ability, respectively.
5.2 Effectiveness on Distillation-Based CL